OpenStudy (anonymous):

I need help solving 3x^3= -24

jhonyy9 (jhonyy9):

like first step divide both sides by 3

OpenStudy (anonymous):

I tried dividing x^3 on both sides, but that didnt make sense

OpenStudy (anonymous):

you divide the notation by 3 too,

OpenStudy (anonymous):

thats the problem

OpenStudy (anonymous):

so now I have x^3=-8, but how do i get rid of exponent?

jhonyy9 (jhonyy9):

will be 3x^3 -24 ----- = ----- 3 3

OpenStudy (anonymous):

so u have x=-8

OpenStudy (anonymous):

cuz u divide ^3 by 3 too so uhave x=-8

OpenStudy (anonymous):

careful alphadog. "so u have x=-8" is incorrect

OpenStudy (anonymous):

ooohhhh sorry

OpenStudy (anonymous):

iget it

OpenStudy (anonymous):

you cannot divide the exponent by 3

OpenStudy (anonymous):

you must take both sides to the ^(1/3) power

OpenStudy (anonymous):

i didnt think so, only the 3's cancel each other out on both sides, correct?

OpenStudy (anonymous):

this makes it x^1 and = -2

OpenStudy (anonymous):

yeah thx, i learnd tht lst yr in algebra, chme thxfor the refresher

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you all, this is a refresher for me as well, havent had to do algebra in years and having problems with my homework